OpenStudy (anonymous):

Find the roots of the equation x^2+ 30 - 13x algebraically

OpenStudy (anonymous):

\begin{aligned} x^2 +30 -13x &= x^2-10x -3x+30 \\ &= x(x-10) -3(x-10) \\ &=(x-10)(x-3) \end{aligned}